# Benchmarking Vision Language Model Unlearning via Fictitious Facial Identity Dataset

### v2

@@ 5 Conclusion
This paper introduces FIUBench, the first unlearning benchmark for Vision Language Models (VLMs) under the Right to be Forgotten setting. After formalizing the VLM unlearning tasks, this benchmark assigns a two-stage evaluation pipeline with our newly proposed Fictitious Facial Identity VQA dataset. Upon applying the unlearning algorithm to the fine-tuned VLM on our dataset, FIUBench offers a comprehensive evaluation by computing both forget quality and model utility, with further robust assessment under membership inference attack and adversarial privacy extraction. Evaluating four baseline unlearning algorithms, our results indicate that none of them achieve good unlearning performance considering both model utility and forget quality. Moreover, the divergent performance of Preference Optimization with and without membership inference attacks underscores the importance of privacy attacks for robust evaluations. We aim to release this benchmark to foster the community’s further research on developing better unlearning methods for VLMs under the setting of Right to be Forgotten.
Limitations. One limitation of our benchmark is that it simulates the Right to be Forgotten scenario in which VLMs are presumed to have rare knowledge by being fine-tuned with our fictitious facial identity knowledge before unlearning. While this approach allows for robust evaluation of unlearning methods, offering consistent and controllable levels of exposed knowledge, it does not account for information acquired during pre-training stages and also in-context learning settings.
@@ D.4 Model Utility-Forget Quality Curve
Figure 4: Model Utility and Forget Quality trade-off curve on LLaVA-Phi-3-mini. Here we convert the Model Utility, EM, and MIA scores to their difference from 100, and then normalize all their values to between 0 and 1. The higher the Forget quality value, the higher the level of forgetfulness of the unlearned models.

### v3

@@ 5 Conclusion
This paper introduces FIUBench, a comprehensive unlearning benchmark for Vision Language Models (VLMs) under the Right to be Forgotten setting. After formalizing the VLM unlearning tasks, this benchmark assigns a two-stage evaluation pipeline with our newly proposed Fictitious Facial Identity VQA dataset. Upon applying the unlearning algorithm to the fine-tuned VLM on our dataset, FIUBench offers a comprehensive evaluation by computing forget quality and model utility, with further robust assessment under membership inference attack and adversarial privacy extraction. Evaluating four baseline unlearning algorithms, our results indicate that none of them achieve good unlearning performance considering both model utility and forget quality. Moreover, the divergent performance of Preference Optimization with and without membership inference attacks underscores the importance of privacy attacks for robust evaluations. We aim to release this benchmark to foster the community’s further research on developing better unlearning methods for VLMs under the setting of Right to be Forgotten.
Limitations
One limitation of our benchmark is that it simulates the Right to be Forgotten scenario in which VLMs are presumed to have rare knowledge by being fine-tuned with our fictitious facial identity knowledge before unlearning. While this approach allows for robust evaluation of unlearning methods, offering consistent and controllable levels of exposed knowledge, it does not account for information acquired during pre-training stages and also in-context learning settings.
Acknowledgements
We sincerely appreciate the reviewers’ insightful comments and valuable feedback, which have greatly enhanced the quality of our manuscript. This work is partially supported by the U.S. Department of Homeland Security under Grant Award Number 17STQAC00001-07-00 and AI2050 program at Schmidt Sciences. Muhao Chen was supported by the DARPA FoundSci Grant HR00112490370 and the NSF of the United States Grant ITE 2333736. Huang is supported by DARPA Transfer from Imprecise and Abstract Models to Autonomous Technologies (TIAMAT) 80321, National Science Foundation NSF-IIS-2147276 FAI, DOD-AFOSR-Air Force Office of Scientific Research under award number FA9550-23-1-0048. Bo Li is supported by the National Science Foundation under Grant No. CNS-2343611, No. 1910100, No. 2046726, NSF AI Institute ACTION No. IIS-2229876, DARPA TIAMAT No. 80321, the National Aeronautics and Space Administration (NASA) under grant No. 80NSSC20M0229, ARL Grant W911NF-23-2-0137, Alfred P. Sloan Fellowship, the research grant from eBay, AI Safety Fund, Virtue AI, U.S. Department of Homeland Security under Grant Award Number 17STQAC00001-07-00 and AI2050 program at Schmidt Sciences.
Disclaimer
The views and conclusions contained in this document are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of the U.S. Department of Homeland Security, the National Science Foundation, DARPA and Schmidt Sciences.
@@ D.4 Model Utility-Forget Quality Curve
Figure 4: Model Utility and Forget Quality trade-off curve on LLaVA-Phi-3-mini. Here we convert the Model Utility, EM, and MIA scores to their difference from 100, then normalize all their values to 0 and 1. The higher the Forget quality value, the higher the level of forgetfulness of the unlearned models.